If a ball is thrown vertically upwards with a velocity of 40m is, then velocity of the ball after two seconds will be (g = 10 m/s2)


        (a)    15 m/s


        (b)    20 m/s


        (c)    25 m/s


        (d)    28 m/s


Grade:10

2 Answers

Jitender Pal
10 years ago

(b)

Initial velocity (u) = 40 m/s

Acceleration a = -g m/s2 = -10 m/s?

Time = 2 seconds

By 1st equation of motion,

v = u + at

v = 40 - 10 (2) = 20 m/s

sharath kumar
10 years ago

v=u+at =>v=40+(-10)2 = 20m/s

option (b)